the preimeter of a rectangle is 43in. more than the length. the width is 10in. find the length of the rectangle
Answer by algebrahouse.com(1659)

"the preimeter of a rectangle is 43in. more than the length. the width is 10in. find the length of the rectangle"

x = length
10 = width
x + 43 = perimeter {perimeter is 43 more than the length}

Perimeter of a rectangle = 2(length) + 2(width)

x + 43 = 2x + 2(10) {substituted into formula for perimeter}
x + 43 = 2x + 20 {multiplied 2 by 10}
43 = x + 20 {subtracted x from both sides}
x = 23 {subtracted 20 from both sides}

23 in. is the length

For more help from me, visit: www.algebrahouse.com